# Self-Supervised Pre-Training with Equilibrium Constraints

### v1

@@ Abstract
Self-supervised pre-training using unlabeled data is widely used in machine learning. In this paper, we propose a new self-supervised pre-training approach to dealing with heterogeneous data. Instead of mixing all the data and minimizing the averaged global loss in the conventional way, we impose additional equilibrium constraints to ensure that the models optimizes each source of heterogeneous data to its local optima after $K$-step gradient descent initialized from the model. We formulate this as a bilevel optimization problem, and use the first-order approximation method to solve the problem. We discuss its connection to model-agnostic meta learning (MAML). Experiments are carried out on self-supervised pre-training using multi-domain and multilingual datasets, demonstrating that the proposed approach can significantly improve the adaptivity of the self-supervised pre-trained model for the downstream supervised fine-tuning tasks.
Index Terms:
self-supervised learning, pre-training, acoustic models, bilevel optimization, automatic speech recogntion
I Introduction
@@
When carrying out self-supervised acoustic pre-training using unlabeled data, it is inevitable to deal with data heterogeneity as the large amount of training data may come from various sources (e.g. domains and languages) bearing different acoustic characteristics. The conventional approach to self-supervised pre-training mixes all data together and minimizes the averaged loss. A drawback of this strategy is that a low averaged global loss does not guarantee a low loss for each source of the heterogeneous data. In this paper, we propose a new self-supervised pre-training approach. In this approach, in addition to the averaged global loss across all heterogeneous data sources, we also impose constraints that require each source of the heterogeneous data also reach its local loss equilibrium after $K$-step gradient descent initialized from the model. Such constraints on local stationarity will ensure that, when optimizing the averaged loss, a good feature representation for each source is also preserved. We formulate this self-supervised pre-training with equilibrium constraints (PTEC) as a bilevel optimization problem [5, 6, 7, 8, 9] where the upper (a.k.a. leader) problem is the conventional averaged global loss across all data sources while the lower (a.k.a. follower) problem is the local loss of each data source. We use the first-order approximation method to solve this bilevel optimization problem and discuss its connection with model-agnostic meta-learning (MAML) [10].
We build pre-trained acoustic models using the proposed PTEC approach and conduct downstream automatic speech recognition (ASR) tasks on two scenarios. One scenario uses speech data from multiple domains and the other uses multilingual speech data. We compare the performance of PTEC with that of the conventional self-supervised pre-trained models. Our experimental results show that the proposed PTEC approach can give rise to a better pre-trained model with superior adaptivity in downstream ASR fine-tuning tasks.
@@
The remainder of the paper is organized as follows. We formulate the problem of PTEC in Section II. Its optimization in given in Section III and the pseudo-code implementation is given in Section IV. The experimental results on multi-domain and multilingual ASR are reported in Section V. Finally, we conclude the paper with a summary in Section VI.
II Problem Formulation
Suppose the unlabeled data $\mathcal{D}$ is collected from $M$ sources
$\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{M}\}.$ (1)
@@ II Problem Formulation
The conventional self-supervised learning, which is henceforth referred to as CSSL, is to train a model that minimizes the following empirical risk
$\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\theta,\mathcal{D}_{i})$ (2)
where $x\in\mathcal{D}_{i}$ is a sample from source $\mathcal{D}_{i}$; $\theta$ is the model parameters; $\ell_{i}(x;\theta,\mathcal{D}_{i})$ is the loss defined on data source $\mathcal{D}_{i}$. In CSSL, the objective function in Eq.2 simply mixes heterogeneous data from various sources together.
@@
Since each source of data may bear its unique characteristics in feature representations, we want the model to preserve these characteristics to offer more robustness and generalization after pre-training. However, this is not guaranteed in Eq.2 by only optimizing towards the averaged global loss. Also considering that the pre-trained model will serve as an initialization point for downstream tasks, we are interested in its adaptivity after multiple steps of gradient descent. To that end, we propose PTEC where equilibrium constraints are imposed in addition to the averaged global loss to ask the model also optimizes each source of heterogeneous data to its local optima after $K$-step gradient descent starting from $\theta$. This is formulated as the following bilevel optimization problem in Eq. 3:
$\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\phi^{*}_{i}(\theta),\mathcal{D}_{i})$
${\rm s.t.}\ \ \ \ \phi^{*}_{i}(\theta)=\operatornamewithlimits{argmin}_{\phi_{i}}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\phi_{i}(\theta),\mathcal{D}_{i}),\ \ i\in[M].$ (3)
@@
where the upper level problem is the averaged global loss with $\theta$ being the initialization model parameter shared by data from all sources and the lower level problem is $M$ local losses with model parameter $\phi_{i}(\theta)$ for each source of data. The dependency of $\phi_{i}(\theta)$ on $\theta$ is through a function of $K$-step gradient descent starting from $\theta$, which will become clear in Section III.
By imposing these equilibrium constraints in the lower level problem, we ensure that the resulting model not only produces good averaged global loss but also serves as a good initialization that can reach a stationary point of the local loss from each data source after a few steps of gradient descent.
III Optimization
@@
To solve the bilevel optimization problem in Eq.3, we first define the following functions to simply the derivation
$f(\theta)$ $\triangleq\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\phi^{*}_{i}(\theta),\mathcal{D}_{i})$ (4)
$g_{i}(\phi_{i})$ $\triangleq\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\phi_{i}(\theta),\mathcal{D}_{i}),\ \ i\in[M]$ (5)
@@
III-A Solving lower level problem
There are $M$ lower level problems defined on each data source. We first solve each of the lower level problems $g_{i}(\phi_{i})$ using K-step gradient descent starting from a common parameter $\theta$ shared by all data sources from the upper level with a learning rate $\alpha$
@@ III-A Solving lower level problem
where $\phi^{i}_{0}=\theta$. $\phi^{i}_{K}$ is used to approximate $\phi^{*}_{i}(\theta)$, a stationary point of $g_{i}(\phi_{i})$:
$\phi^{*}_{i}(\theta)\approx\phi^{i}_{K}.$ (7)
Particularly, when $K=1$, we have
@@
The dependency of $\phi_{i}(\theta)$ on $\theta$ is obvious in Eq. 6 and Eq.8 as a special case when $K=1$.
III-B Solving upper level problem
The upper level problem $f(\theta)$ is the global loss of all data sources using a shared parameter $\theta$. We solve it also using gradient descent with a learning rate $\beta$
@@ III-B Solving upper level problem
$\nabla_{\theta}f(\theta)$ $=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}\left(\sum_{x\in\mathcal{D}_{i}}\ell_{i}(x;\phi^{*}_{i}(\theta),\mathcal{D}_{i})\right)$
$=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{*}_{i}(\theta))$
$\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{i}_{K})$ (10)
The last step is due to Eq.7 where $\phi^{*}_{i}(\theta)$ is approximated by $\phi^{i}_{K}$.
@@
$\nabla_{\theta}f(\theta)$ $\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{i}_{K})$
$=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})\nabla_{\theta}(\phi^{i}_{K})$
@@
In the last step
$\nabla_{\theta}(\phi^{i}_{0})=\mathbf{I}$ (12)
which is due to
$\phi^{i}_{0}=\theta$ (13)
since the gradient descent starts from the shared parameter $\theta$.
We then expand each $\phi^{i}_{k}$ with its gradient descent update
$\nabla_{\theta}f(\theta)$ $\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})\prod_{k=1}^{K}\nabla_{\phi^{i}_{k-1}}(\phi^{i}_{k})$
$=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})\prod_{k=1}^{K}\nabla_{\phi^{i}_{k-1}}[\phi^{i}_{k-1}-\alpha\nabla_{\phi_{i}}g_{i}(\phi^{i}_{k-1})]$
@@
Particularly, if we only conduct gradient descent on the lower level for just one step (i.e. $K=1$), we have
$\nabla_{\theta}f(\theta)\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{1})$ (16)
IV Implementation
Based on the derivation in Secs. III-A and III-B, the pseudo-code implementation of PTEC is given in Algorithm 1.
@@ IV Implementation
   Update global model $\theta_{t}\leftarrow\theta_{t-1}-\beta\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})$;
@@
Algorithm 1 Self-Supervised Pre-Training with Equilibrium Constraints (PTEC)
@@
As shown in Algorithm 1, the PTEC training is carried out in $T$ iterations where the problems of the lower and upper levels are alternately optimized. Specifically, in each iteration, the $M$ lower level problems on each data source are first (approximately) solved individually using $K$-step gradient descent. The gradient descent is initialized with the global model parameter from the upper level which is the same for each of the $M$ problems. After the lower level problem is (approximately) solved, the resulting local optimum of each problem is used to evaluate its gradient. The upper level problem is then optimized using gradient descent based on the averaged gradients evaluated from the $M$ lower level problems. Note that the local updates on $M$ sources can be performed in parallel. We find that in order for PTEC to perform well it needs to be appropriately initialized. In our experiments, we always initialize PTEC with a model trained from CSSL. Based on the observation in [14], when a deep model is sufficiently trained, the majority of its eigenvalues of the Hessian matrix of the loss function tends to be zeros, which indicates a flat curvature. This would make the assumption in Eq.15 more accurate and therefore the first-order approximation more legitimate.
@@
When considering the first-order approximation with one step gradient descent, the proposed PTEC shares similarity with MAML [10] from the optimization perspective as MAML can also be interpreted from a bilevel optimization framework [15, 16]. There are works using MAML in acoustic modeling [17, 18, 19, 20] most of which are on supervised learning. However, there are differences between the two. The proposed equilibrium-constrained pre-training is on self-supervised learning without relying on ground-truth labels. Furthermore, it uses the same training data in both the upper and lower level problems without any validation data involved. Therefore, there is no so-called “meta-train” and “meta-test” analogy drawn in the design [10, 16]. The similarity between the two is the outcome of the first-order approximation under gradient unrolling to the solution of a bilevel optimization problem.
V Experiments
@@
We evaluate the proposed PTEC on two sets of experiments. One is self-supervised English acoustic model pre-training using speech data from multiple domains based on BEST-RQ [4]. The other is self-supervised multilingual acoustic model pre-training based on contrastive predictive coding (CPC) [21].
V-A Multi-Domain Pre-Training
@@
In this experiment, we use English speech data collected from a broad variety of domains. All the speech data has a sampling rate of 16KHz. Table I gives the details of domain distribution and the amounts of data from each domain. Data is collected from a total of nine domains including Broadcast News, ViaVoice dictation, AMI meetings, Great Britain (GB) English, Librispeech, Australian (AU) English, Hospitality and Accented English (Asian and Latin). Among these nine domains, data from the first five domains is used for pre-training while that from the last five is used for downstream fine-tuning and test. Therefore, in the downstream fine-tuning tasks, two domains are seen in the training (GB English and Librispeech) and three domains are unseen from the training (AU English, Hospitality and Accented). The data is also unbalanced in amount where some domain can have as many as 860 hours of speech while the others can have only 40 hours. We use this setting to simulate the real-world application scenarios. The hours of speech on test sets are also shown in the table. Note that Librispeech (test-clean/test-other) and Accented (asian/latin) both have two test sets by themselves.
@@
TABLE I: Speech data from multiple domains used in self-supervised pre-training and downstream fine-tuning/test tasks.
The pre-trained acoustic model is a Conformer model [22]. The input is 40-dimensional log-Mel spectrogram features and their first and second-order derivatives. Features of every two adjacent frames are concatenated which gives 240-dimensional input vectors. The model has 10 conformer blocks. Each block has 512 hidden units and 8 attention heads of 64 dimensions. The convolution kernel size is 31. The total number of parameters is 70M. The self-supervised training is carried out using the BEST-RQ criterion. The masking probability in BEST-RQ is 0.02. The mask span is 20 frames. The masked frames are replaced with Gaussian noise with 0 mean and 0.1 variance. The size of the random codebook is 256.
When training the CSSL as the baseline, we start the training with a learning rate 2e42e\!-\!4 for 60 epochs which is then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward. The training ends after 80 epochs. When training the model using PTEC, we start the training with a local learning rate $\alpha=1e\!-\!4$ and a global learning rate $\beta=1e\!-\!5$ for 40 epochs. They are then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward synchronously. The training ends after 60 epochs. In PTEC, the lower level optimization on each data source is conducted in parallel. To deal with the unbalanced data size across different data sources, we make the batch size roughly proportional to the total amount of data to make sure each data source produces about the same number of batches. In addition, we also make random skipping of batches during the training when a data source has more batches than the others. All training uses the AdamW optimizer.
After pre-training, a linear layer is added to the pre-trained conformer model. It is then fine-tuned with labeled speech data on each of the downstream ASR tasks using the Connectionist Temporal Classification (CTC) [23] criterion. The softmax layer contains 43 output units which correspond to 42 characters and the null symbol. The fine-tuning starts with a learning rate $\beta=1e\!-\!4$ for 5 epochs which is then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward. The fine-tuning ends after 15 epochs.
Table II compares word error rates (WERs) of CSSL as the baseline and the proposed PTEC on five downstream ASR tasks. Each task represents an ASR application in a particular domain. We also compare the performance of PTEC with different $K$-step local updates in the lower level optimization. It can be observed from the table that PTEC shows improvements over CSSL on seven test sets across all downstream tasks. It also can be observed that running more local updates (KK=3) may not necessarily yield better performance but increases the computational cost. Therefore, we will stick to KK=1 in the remaining experiments.
@@
PTEC (KK=1) 18.5 15.8/17.0 24.6 24.0 11.5/22.7
PTEC (KK=3) 18.7 15.7/17.1 26.9 24.5 11.5/22.7
TABLE II: WERs of CSSL and PTEC on five downstream ASR tasks.
In the experiments in Table II we use CSSL to initialize PTEC. This procedure can be iterative where we can use CSSL and PTEC to perform mutual initialization. By doing this, both landscapes of the upper global loss and lower local loss will be further explored and hopefully we can end up with a better optimum. The results are given in Table III. We conduct three rounds of CSSL and PTEC sequentially (denoted CSSL.ii and PTEC.ii, respectively, $i=1,2,3$). We initialize PTEC.ii using CSSL.ii and initialize CSSL.i+1i\!+\!1 using PTEC.ii, In every round, CSSL runs 80 epochs and PTEC runs 60 epochs, which is the same training schedule as that in Table II. The results clearly show that this iterative training strategy can greatly improve the adaptivity of the pre-trained model which obtains significant WER reduction in all downstream ASR tasks. The results also show that CSSL does not always improve over the PTEC model which initializes it on all downstream domains. This is because the averaged global loss can not guarantee good performance on each specific domain. On the other hand, PTEC always outperforms its initial CSSL model in every downstream domain and hence demonstrates its superior robustness. Overall, if we compare the final WERs after the iterative training (the last row) to the baseline (the first row), that gives rise to 15%-40% relative improvements across the seven test sets and the improvements are consistent.
Note that the WERs of the Librispeech baseline (12.4/24.5) use only 100 hours of labeled data for supervised fine-tuning. It should not be confused with WERs using 960 hours of labeled data commonly reported in literature. Furthermore, its data distribution is further flattened by data from various domains. (As a reference, if we only use 860 hours of unlabeled speech and remove data from other domains in CSSL.1, the WERs are 7.4/20.1). After PTEC training, the WERs are improved to 9.4/18.2, which are competitive, especially for the mismatched case in test-other [24]. This is achieved without sacrificing the performance of other test sets.
@@
PTEC.3 15.6 13.7/15.1 21.3 18.4 9.4/18.2
TABLE III: WERs of iterative CSSL and PTEC on five downstream fine-tuning ASR tasks. In the training, CSSL and PTEC alternatingly initialize each other.
V-B Multilingual Pre-Training
For the multilingual pre-training experiment, we use a subset of the multilingual data from the CoVoST v2 dataset [25], which is sampled at 48 kHz. Table IV provides details on the language distribution and the amount of data for each language. The data spans seven languages: English, French, Dutch, Turkish, Swedish, Tamil, and Welsh. The first three languages (English, French, Dutch) are used for pre-training and the remaining four (Turkish, Swedish, Tamil, Welsh) are used for downstream fine-tuning ASR tasks. Unlike the multi-domain experiment in Sec.V-A, there is no overlap between the pre-training and fine-tuning languages in this experiment. However, similar to the multi-domain experiment, both the pre-training and fine-tuning data are unbalanced in quantity. For instance, we use a 357-hour subset of the English data for pre-training, whereas some languages have as little as one hour of data.

### v2

# Heterogeneous Self-Supervised Acoustic Pre-Training with Local Constraints

@@ Abstract
Self-supervised pre-training using unlabeled data is widely used in automatic speech recognition. In this paper, we propose a new self-supervised pre-training approach to dealing with heterogeneous data. Instead of mixing all the data and minimizing the averaged global loss in the conventional way, we impose additional local constraints to ensure that the model optimizes each source of heterogeneous data to its local optimum after $K$-step gradient descent initialized from the model. We formulate this as a bilevel optimization problem, and use the first-order approximation method to solve the problem. We discuss its connection to model-agnostic meta learning. Experiments are carried out on self-supervised pre-training using multi-domain and multilingual datasets, demonstrating that the proposed approach can significantly improve the adaptivity of the self-supervised pre-trained model for the downstream supervised fine-tuning tasks.
Index Terms—  self-supervised learning, pre-training, acoustic models, bilevel optimization, automatic speech recognition
1 Introduction
@@
When carrying out self-supervised acoustic pre-training using unlabeled data, it is inevitable to deal with data heterogeneity as the large amount of training data may come from various sources (e.g., domains and languages) bearing different acoustic characteristics. The conventional approach to self-supervised pre-training mixes all data together and minimizes the averaged loss. A drawback of this strategy is that a low average global loss does not guarantee a low loss for each source of the heterogeneous data. In this paper, we propose a new self-supervised pre-training approach. In this approach, in addition to the averaged global loss across all heterogeneous data sources, we also impose constraints that require each source of the heterogeneous data to reach its local loss optimum after $K$-step gradient descent initialized from the model. Such constraints on the local optimum will ensure that, when optimizing the averaged loss, a good feature representation for each source is also preserved. We formulate this self-supervised pre-training with local constraints (PTLOC) as a bilevel optimization problem [5, 6, 7, 8, 9] where the upper problem is the averaged global loss across all data sources, while the lower problem is the local loss of each data source. We use the first-order approximation method to solve this bilevel optimization problem and discuss its connection with model-agnostic meta-learning (MAML) [10]. We build pre-trained acoustic models using the proposed PTLOC approach and conduct downstream automatic speech recognition (ASR) tasks on two scenarios. One scenario uses speech data from multiple domains and the other uses multilingual speech data. We compare the performance of PTLOC with that of the conventional self-supervised pre-trained models. Our experiments show that the proposed PTLOC can give rise to a better pre-trained model with superior adaptivity in downstream ASR fine-tuning (FT) tasks.
@@
The remainder of the paper is organized as follows. We formulate the problem of PTLOC in Section 2. Its optimization in given in Section 3 and the pseudo-code implementation is given in Section 4. The experimental results on multi-domain and multilingual ASR are reported in Section 5. Finally, we conclude the paper with a summary in Section 6.
2 Problem Formulation
@@
Suppose the unlabeled data $\mathcal{D}$ is collected from $M$ sources: $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{M}\}$. Conventional self-supervised learning (CSSL) trains a model that minimizes the following empirical risk
$\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\theta;x,\mathcal{D}_{i})$ (1)
@@
where $x\in\mathcal{D}_{i}$ is a sample from source $\mathcal{D}_{i}$; $\theta$ is the model parameters; $\ell_{i}(\theta;x,\mathcal{D}_{i})$ is the loss defined on data source $\mathcal{D}_{i}$. In CSSL, the objective function in Eq. (1) simply mixes heterogeneous data from various sources together. Since each source of data may bear its unique characteristics in feature representations, we want the model to preserve these characteristics to offer more robustness and generalization after pre-training. However, this is not guaranteed in Eq. (1) by only optimizing towards the averaged global loss. Also considering that the pre-trained model will serve as an initialization point for downstream tasks, we are interested in its adaptivity after multiple steps of gradient descent. To that end, we propose PTLOC, where local constraints for each data source are imposed in addition to the global averaged loss, requiring that the model also optimize each source of heterogeneous data to its local optimum after $K$-step gradient descent initialized from $\theta$. This is formulated as the following bilevel optimization problem:
$\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\phi^{*}_{i}(\theta);x,\mathcal{D}_{i})$
${\rm s.t.}\ \ \ \ \phi^{*}_{i}(\theta)=\operatornamewithlimits{argmin}_{\phi_{i}}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\phi_{i}(\theta);x,\mathcal{D}_{i}),\ \ i\in[M].$ (2)
@@
where the upper level problem is the averaged global loss with $\theta$ being the initialization model parameter shared by data from all sources and the lower level problem is $M$ local losses with model parameter $\phi_{i}(\theta)$ for each data source. The dependency of $\phi_{i}(\theta)$ on $\theta$ is through a function of $K$-step gradient descent starting from $\theta$, which will become clear in Section 3. By imposing the constraints in the lower-level problem, we ensure that the resulting model not only produces a good average global loss but also serves as a good initialization that can reach a local optimal point of the local loss from each data source after a few steps of gradient descent.
3 Optimization
@@
To solve the bilevel optimization problem in Eq. (2), we first define the following functions to simplify the derivation
$f(\theta)$ $\triangleq\frac{1}{M}\sum_{i=1}^{M}\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\phi^{*}_{i}(\theta);x,\mathcal{D}_{i})$ (3)
$g_{i}(\phi_{i})$ $\triangleq\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\phi_{i}(\theta);x,\mathcal{D}_{i}),\ \ i\in[M]$ (4)
@@
3.1 Solving the lower level problem
There are $M$ lower-level problems, one for each data source. We first solve each of the lower-level problems $g_{i}(\phi_{i})$ using K-step gradient descent starting from a common parameter $\theta$ shared by all data sources from the upper level, with a learning rate $\alpha$
@@ 3.1 Solving the lower level problem
where $\phi^{i}_{0}=\theta$. $\phi^{i}_{K}$ is used to approximate $\phi^{*}_{i}(\theta)$: $\phi^{*}_{i}(\theta)\approx\phi^{i}_{K}$. Particularly, when $K=1$, we have
@@
The dependency of $\phi_{i}(\theta)$ on $\theta$ is obvious in Eq. 5 with Eq. (6) being a special case when $K=1$.
3.2 Solving the upper level problem
The upper level problem $f(\theta)$ is the global loss of all data sources starting from a shared parameter $\theta$. We also solve it using gradient descent with a learning rate $\beta$
@@ 3.2 Solving the upper level problem
$\nabla_{\theta}f(\theta)$ $=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}\left(\sum_{x\in\mathcal{D}_{i}}\ell_{i}(\phi^{*}_{i}(\theta);x,\mathcal{D}_{i})\right)$
$=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{*}_{i}(\theta))\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{i}_{K})$ (8)
The last step is because $\phi^{*}_{i}(\theta)$ is approximated by $\phi^{i}_{K}$.
@@
$\nabla_{\theta}f(\theta)$ $\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\theta}g_{i}(\phi^{i}_{K})=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})\nabla_{\theta}(\phi^{i}_{K})$
@@
In the last step $\nabla_{\theta}(\phi^{i}_{0})\!=\!\mathbf{I}$ which is due to $\phi^{i}_{0}\!=\!\theta$ since the gradient descent starts from the shared parameter $\theta$.
We then expand each $\phi^{i}_{k}$ with its gradient descent update in Eq. (5)
$\nabla_{\theta}f(\theta)$ $\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})\prod_{k=1}^{K}\nabla_{\phi^{i}_{k-1}}[\phi^{i}_{k-1}-\alpha\nabla_{\phi_{i}}g_{i}(\phi^{i}_{k-1})]$
@@
Particularly, if we only conduct gradient descent on the lower level for just one step (i.e., $K=1$), we have
$\nabla_{\theta}f(\theta)\approx\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{1}).$ (12)
4 Implementation
@@
  Compute global gradient $\nabla_{\theta}f(\theta_{t-1})=\frac{1}{M}\sum_{i=1}^{M}\nabla_{\phi_{i}}g_{i}(\phi^{i}_{K})$;
 Update global model
     $\theta_{t}\leftarrow\theta_{t-1}-\beta\nabla_{\theta}f(\theta_{t-1})$;
@@
Algorithm 1 Self-Supervised Pre-Training with Local Constraints (PTLOC)
@@
Based on the derivation in Secs. 3.1 and 3.2, the pseudo-code implementation of PTLOC is given in Algorithm 1. The PTLOC training is carried out in $T$ iterations, where the problems of the lower and upper levels are alternately optimized. Specifically, in each iteration, the $M$ lower-level problems on each data source are first (approximately) solved individually using $K$-step gradient descent. The gradient descent is initialized with the global model parameter from the upper level, which is the same for each of the $M$ problems. After the lower-level problem is (approximately) solved, the resulting local optimum of each problem is used to evaluate its gradient. The upper-level problem is then optimized using gradient descent based on the averaged gradients evaluated from the $M$ lower-level problems. We find that in order for PTLOC to perform well, it needs to be appropriately initialized. In our experiments, we always initialize PTLOC with a model trained from CSSL. Based on the observation in [14], when a deep model is sufficiently trained, the majority of the eigenvalues of the Hessian matrix of the loss function tend to be zero. This makes the assumption in Eq. (11) more accurate and therefore the first-order approximation more legitimate.
@@
When considering the first-order approximation with one step gradient descent, PTLOC shares similarity with MAML from the optimization perspective, as MAML can also be interpreted from a bilevel optimization framework [15, 16]. There are works using MAML in acoustic modeling [17, 18, 19, 20], most of which are on supervised learning. However, there are differences between the two: i) PTLOC is on self-supervised learning without relying on ground-truth labels; and, ii) PTLOC uses the same training data in both the upper and lower level problems without any validation data as those in the meta learning design [10, 16]. The similarity between the two is the outcome of the first-order approximation to the solution of a bilevel optimization problem.
5 Experiments
@@
We evaluate the proposed PTLOC on two sets of experiments. One is a self-supervised English acoustic model pre-training using speech data from multiple domains based on BEST-RQ [4]. The other is a self-supervised multilingual acoustic model pre-training based on contrastive predictive coding (CPC) [21].
5.1 Multi-Domain Pre-Training
@@
In this experiment, we use English data collected from a broad variety of domains. All the speech signals have a sampling rate of 16KHz. Table 1 gives the details of the domain distribution and the amount of data from each domain. Data is collected from a total of eight domains, including Broadcast News, ViaVoice dictation, AMI meetings, British (GB) English, Librispeech, Australian (AU) English, Hospitality, and Accented English (Asian and Latin). Among these eight domains, data from the first five domains is used for pre-training, while that from the last five is used for downstream FT and test. Therefore, in the downstream FT tasks, two domains are seen in the pre-training, and three domains are unseen. The data is unbalanced in amount. We use this setting to simulate real-world application scenarios. The hours of speech on test sets are shown in the table. Note that Librispeech (clean/other) and Accented (Asian/Latin) both have two test sets.
@@
Table 1: Speech data from multiple domains used in self-supervised pre-training and downstream fine-tuning/test tasks.
The pre-trained acoustic model is a Conformer model [22]. The input is 40-dimensional log-Mel spectrogram features and their first and second-order derivatives. Features of every two adjacent frames are concatenated which gives 240-dimensional input vectors. The model has 10 conformer blocks. Each block has 512 hidden units and 8 attention heads of 64 dimensions. The convolution kernel size is 31. The total number of parameters is 70M. The self-supervised training is carried out using BEST-RQ. The masking probability in BEST-RQ is 0.02. The mask span is 20 frames. The masked frames are replaced with Gaussian noise with 0 mean and 0.1 variance. The size of the random codebook is 256.
For the CSSL baseline, we start the training with a learning rate 2e42e\!-\!4 for 60 epochs which is then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward. The training ends after 80 epochs. For PTLOC, we start the training with a local learning rate $\alpha=1e\!-\!4$ and a global learning rate $\beta=1e\!-\!5$ for 40 epochs. They are then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward synchronously. The training ends after 60 epochs. The lower-level optimization on each data source is conducted in parallel. To deal with the unbalanced data size across different data sources, we make the batch size roughly proportional to the total amount of data to make each data source produce about the same number of batches. In addition, we also make random skipping of batches during the training when a data source has more batches than the others. All training uses the AdamW optimizer.
After pre-training, a linear layer is added to the pre-trained conformer model for FT with labeled data on each of the downstream ASR tasks using the Connectionist Temporal Classification (CTC) [23] criterion. The softmax layer contains 43 output units, including 42 characters and the null symbol. The FT starts with a learning rate $\beta=1e\!-\!4$ for 5 epochs which is then annealed by $\frac{1}{\sqrt{2}}$ every epoch afterward. The FT ends after 15 epochs.
@@
PTLOC (KK=1) 18.5 15.8/17.0 24.6 24.0 11.5/22.7
PTLOC (KK=3) 18.7 15.7/17.1 26.9 24.5 11.5/22.7
Table 2: WERs of CSSL and PTLOC on five downstream ASR tasks with different $K$.
Table 2 compares word error rates (WERs) of CSSL as the baseline and the proposed PTLOC on five downstream ASR tasks. Each task represents an ASR application in a particular domain. We also compare the performance of PTLOC with different $K$-step local updates in the lower-level optimization. It can be observed that PTLOC shows improvements over CSSL on seven test sets across all downstream tasks. It can also be observed that running more local updates (KK=3) may not necessarily yield better performance, but increases the computational cost. Therefore, we will stick to KK=1 in the remaining experiments.
In the experiments in Table 2, we use CSSL to initialize PTLOC. This procedure can be iterative where we can use CSSL and PTLOC to perform mutual initialization. By doing this, both landscapes of the upper global loss and lower local loss will be further explored and hopefully we can end up with a better optimum. The results are given in Table 3. We conduct three rounds of CSSL and PTLOC sequentially (denoted CSSL.ii and PTLOC.ii, respectively, $i=1,2,3$). We initialize PTLOC.ii using CSSL.ii and initialize CSSL.i+1i\!+\!1 using PTLOC.ii. In every round, we use the same training schedule as that in Table 2. The results clearly show that this iterative training strategy can greatly improve the adaptivity of the pre-trained model which obtains significant WER reduction in all downstream ASR tasks. The results also show that CSSL does not always improve over the PTLOC model which initializes it on all downstream domains. This is because the averaged global loss can not guarantee good performance on each specific domain. On the other hand, PTLOC always outperforms its initial CSSL model in every downstream domain and hence demonstrates its superior robustness. Overall, if we compare the final WERs after the iterative training (the last row) to the baseline (the first row), that gives rise to 15%-40% relative improvements across the seven test sets and the improvements are consistent. Note that the WERs of the Librispeech baseline (12.4/24.5) use only 100 hours of labeled data for supervised FT. It should not be confused with WERs using 960 hours of labeled data, commonly reported in the literature. Furthermore, its distribution is further flattened by data from various domains. (As a reference, if we only use 860 hours of unlabeled speech and remove data from other domains in CSSL.1, the WERs are 7.4/20.1).
@@
PTLOC.3 15.6 13.7/15.1 21.3 18.4 9.4/18.2
Table 3: WERs of iterative CSSL and PTLOC on five downstream fine-tuning ASR tasks. In the training, CSSL and PTEC alternately initialize each other.
5.2 Multilingual Pre-Training
For the multilingual pre-training experiment, we use a subset of the multilingual data from the CoVoST v2 dataset [24], which is sampled at 48 kHz. Table 4 provides details on the language distribution and the amount of data for each language. The data includes seven languages. The first three languages (English, French, Dutch) are used for pre-training and the remaining four (Turkish, Swedish, Tamil, Welsh) are used for downstream FT ASR tasks. Unlike the multi-domain experiment in Sec.5.1, there is no overlap between the pre-training and FT languages in this experiment. But both the pre-training and FT data are unbalanced in quantity.